Find the soultion to the system of equations.
x + 7y = -2
3x + y = 34

Answer by checkley77(12569)

x + 7y = -2 multiply this equation by -3 & add.
3x + y = 34
-3x-21y=6
-------------------
-20y=40
y=40/-20
y=-2 ans.
x-7*-2=-2
x-14=-2
x=14-2
x=12 ans.
Proof:
3*12-2=34
36-2=34
34=34